Place the following set of numbers in ascending order.
-2/5,1/6, 7/8, 5/9, 3/4

Answer by carter6949(26)

ok the def. of assinding order is from least to gratest
so look at the numbers you have.. -2/5, 1/6, 7/8, 5/9, 3/4
you could divide the numbers (-2/5 = .4), (1/6 = .16), (7/8 = .87), (5/9 = .56), (3/4 = .75)
which is the least? -2/5, 1/6, 5/9, 3/4, 7/8
next least